# Regularization of the ensemble Kalman filter using a non-parametric, non-stationary spatial model

### v3

@@ Appendix I Consistency of the estimator of local spectra
The three parameter fields are at the second level in the hierarchy. $s(x)$ equals the spatially variable standard deviation of the random field in question. $\lambda(x)$ is the local length scale parameter. $\gamma(x)$ is the local shape parameter. These three fields are set to be non-Gaussian stationary random fields computed as follows:
@@
Here $g$ is the nonlinear transformation function defined in the next paragraph and $\chi_{S},\chi_{\lambda},\chi_{\gamma}$ are the three pre-transform fields (introduced in the next subsection). The pre-transform fields are independent stationary zero-mean Gaussian fields. The hyperparameters with the subscript ${}_{\rm add}$ prohibit unrealistically small values in the respective parameter fields and, together with the hyperparameters with the subscript ${}_{\rm mult}$, specify the median values of the respective parameter fields. The hyperparameters $\varkappa_{\bullet}>0$ affect the magnitude of deviations of the parameter fields from their median values, causing spatial non-stationarity in the local spectra.
The transformation function $g$ in Eq. (82) is taken, following (Tsyrulnikov and Rakitko, 2019), to be the scaled and shifted logistic function:
@@
Due to the nonlinearity of the transformation function $g$, the above parameter fields $s(x),\lambda(x),\gamma(x)$ are non-Gaussian. Their point-wise distributions are known as logit-normal or logit-Gaussian. If some $\varkappa_{\bullet}=1$, the respective parameter field is constant in space. The higher $\varkappa_{\bullet}$, the more variable in space becomes the respective parameter. In the numerical experiments described below, all $\varkappa_{\bullet}$ are the same, equal to $\varkappa$. We refer to $\varkappa$ as the non-stationarity strength hyperparameter (as it determines $S$, see Eq. (55)).
J.3 Pre-transform fields
@@
where
$\Lambda_{\rm NSL}=(\lambda_{\rm add}+{\lambda}_{\rm mult})\cdot\mu_{\rm NSL}$ (85)
is the non-stationarity length scale of $\xi(x)$, $\mu_{\rm NSL}$ is the non-stationarity length scale hyperparameter (cf. G), and $\Gamma=\gamma_{\rm add}+\gamma_{\rm mult}$.
Back substitution of the pre-transform random fields $\chi_{S},\chi_{\lambda},\chi_{\gamma}$ into Eq. (82) followed by plugging in the resulting parameter fields $s(x)$, $\lambda(x)$, and $\gamma(x)$ into Eq. (81) (after $c(x)$ is adjusted point-wise as explained in J.1) yields the non-stationary random local spectra $f_{\ell}(x)$. Given the local spectra, the random field $\xi$ is computed as explained above in J.1.
J.4 Hyperparameters
@@
In the experiments reported in section 5, two hyperparameters were varied: $\varkappa$ and $\mu_{\rm NSL}$. We recall that $\varkappa$ determines the strength of the spatial non-stationarity whereas $\mu_{\rm NSL}$ specifies the length scale of the non-stationarity pattern. To demonstrate how the non-stationarity length scale parameter $\mu_{\rm NSL}$ influences the structure of the random field $\xi$, we computed666 We used SHTools (Wieczorek and Meschede, 2018) to generate pseudo-random fields on the sphere. two realizations of $\xi$, one generated with the default $\mu_{\rm NSL}=3$ (see the right panel of Fig. 9) and the other with the smaller $\mu_{\rm NSL}=1$ (left panel of Fig. 9). In both panels of Fig. 9, one can see a substantial degree of spatial non-stationarity, both in the field’s magnitude and length scale. With $\mu_{\rm NSL}=1$ (Fig. 9(left)), areas with a small length scale of $\xi$ (seen as patchy spots) and with a large length scale (smooth spots) are both seen to be smaller in size compared with those in the right panel. This reflects a smaller non-stationarity length scale with $\mu_{\rm NSL}=1$ than with $\mu_{\rm NSL}=3$, as expected.
Figure 9: Two realizations of the random field produced by the parametric model of “truth”. Left: $\mu_{\rm NSL}=1$. Right: $\mu_{\rm NSL}=3$. x-axis: longitude index, y-axis: co-latitude index.
J.5 Local stationarity
The (random) local spectra $f_{\ell}$ defined above in this Appendix give rise to a locally stationary random field $\xi(x)$ if $\varkappa\approx 1$ (low non-stationarity strength) or $\mu_{\rm NSL}\gg 1$ (large non-stationarity length), cf. G.
@@
If $\mu_{\rm NSL}\gg 1$, the non-stationarity length $\Lambda_{\rm NSL}$ (Eq. (85)) is much larger than the median local length scale $\lambda_{\rm add}+\lambda_{\rm mult}$ implying local stationarity according to G.
Appendix K Details on the neural network
@@
In the training process, we used the Adam optimizer (a stochastic gradient-based numerical optimization algorithm, e.g., Kingma and Ba (2014), Goodfellow et al. (2016)). The network’s hyperparameters were taken according to recommendations in Kingma and Ba (2014). The size of the minibatch (a random subsample without replacement of the training sample, used to compute an estimate of the current gradient of the loss function with respect to the network weights) was tuned to be 2500. The number of iterations over the whole training sample (epochs) was 30 on the sphere and 200 on the circle.

### v5

@@ H.2 Parameter fields
The random field $s(x)$ equals the spatially variable standard deviation of the random field in question. $\lambda(x)$ is the local length scale parameter. $\gamma(x)$ is the local shape parameter. These three fields are set to be non-Gaussian stationary random fields computed as follows:
@@
Here $g$ is the nonlinear transformation function defined in the next paragraph and $\chi_{S},\chi_{\lambda},\chi_{\gamma}$ are the three pre-transform fields (introduced in the next subsection). The pre-transform fields are independent stationary zero-mean Gaussian fields. The hyperparameters with the subscript add prohibit unrealistically small values in the respective parameter fields and, together with the hyperparameters with the subscript mult, specify the median values of the respective parameter fields. The hyperparameters $\varkappa_{\bullet}>0$ affect the magnitude of deviations of the parameter fields from their median values.
The transformation function $g$ in Eq. (84) is taken, following Tsyrulnikov and Rakitko (2019), to be the scaled and shifted logistic function:
@@
Due to the nonlinearity of the transformation function $g$, the above parameter fields $s(x),\lambda(x),\gamma(x)$ are non-Gaussian. Their point-wise distributions are known as logit-normal or logit-Gaussian. If some $\varkappa_{\bullet}=1$, the respective parameter field is constant in space. The higher $\varkappa_{\bullet}$, the more variable in space becomes the respective parameter. In the numerical experiments described below, all $\varkappa_{\bullet}$ are the same, equal to $\varkappa$. We refer to $\varkappa$ as the non-stationarity strength hyperparameter because it specifies the degree of spatial deviations of the fields $s(x)$, $\lambda(x)$, and $\gamma(x)$ (according to Eq. (84)) and thus determines the magnitude of spatial variability in the spectral functions $\sigma_{\ell}(x)$, see Eq. (83), which cause non-stationarity.
H.3 Pre-transform fields
@@
where $\Gamma=\gamma_{\rm add}+\gamma_{\rm mult}$,
$\Lambda_{\rm NSL}=L_{\rm med}\cdot\mu_{\rm NSL}$ (87)
is the non-stationarity length scale,
$L_{\rm med}=\lambda_{\rm add}+{\lambda}_{\rm mult}$ (88)
is the median process length scale, and $\mu_{\rm NSL}$ is the non-stationarity length scale hyperparameter.
Back substitution of the pre-transform random fields $\chi_{S},\chi_{\lambda},\chi_{\gamma}$ into Eq. (84) followed by plugging in the resulting parameter fields $s(x)$, $\lambda(x)$, and $\gamma(x)$ into Eq. (83) (after $c(x)$ is adjusted point-wise as explained in H.1) yields the non-stationary random spectral functions $\sigma_{\ell}(x)$. Given $\sigma_{\ell}(\cdot)$, the random field $\xi(\cdot)$ is computed using Eq. (10).
H.4 Hyperparameters
@@
H.5 Local stationarity
The (random) local spectra $f_{\ell}$ defined above in this Appendix give rise to a locally stationary random field $\xi(x)$ if $\varkappa\approx 1$ (low non-stationarity strength) or $\mu_{\rm NSL}\gg 1$ (large non-stationarity length), cf. section 3.6.
@@ H.5 Local stationarity
If $\mu_{\rm NSL}\gg 1$, the non-stationarity length $\Lambda_{\rm NSL}$ (Eq. (87)) is much larger than the median local length scale $L_{\rm med}$, implying local stationarity according to section 3.6.
Appendix I Details on the neural network
@@
In the training process, we used the Adam optimizer (a stochastic gradient-based numerical optimization algorithm (e.g. Kingma and Ba, 2014, Goodfellow et al., 2016). The network’s hyperparameters were taken according to recommendations in Kingma and Ba (2014). The size of the minibatch (a random subsample without replacement of the training sample, used to compute an estimate of the current gradient of the loss function with respect to the network weights) was tuned to be 2500. The number of iterations over the whole training sample (epochs) was 30 on the sphere and 200 on the circle.